# What is the Plausible Memory Representation?

### v1

@@ 4.2 Do TG’s and Transformer’s NAE have independent contributions?
4.3 What aspect of memory retrieval do TG’s and Transformer’s NAE capture?
To investigate the aspects of human memory retrieval captured by TG’s and Transformer’s NAE, we analyzed differences in prediction improvement across part-of-speech (POS) tags annotated in the Natural Stories corpus. Our analysis followed three steps: (i) selecting POS tags with more than 1,000 occurrences, (ii) for each POS tag, testing the significance of improvement from the baseline regression model (measured in Root Mean Squared Error, RMSE) when adding NAE of the current and previous word as fixed effects,111111We used the same regression models as in Section 4.2, where surprisal and NAE values were averaged across seeds. and (iii) examining the significance of differences in RMSE between TG and Transformer for POS tags where either model showed significant improvement. We assessed significance using Wilcoxon signed-rank tests with Bonferroni correction ().
Figure 3 presents the differences in prediction improvement across POS tags.121212For a complete list of POS tags in the Natural Stories corpus, see Appendix C. Consistent with the larger LogLik value, TG’s NAE demonstrated advantages over Transformer’s NAE across a broader range of POS tags. Notably, TG’s NAE exhibited superior improvement across verbs (VB, VBD, VBG, VBP), while Transformer’s NAE excelled only for possessive pronouns ($PRP). These findings indicate that different types of retrieval operations—verb-triggered retrieval (e.g., subject retrieval) and possessive pronoun-triggered retrieval (e.g., antecedent retrieval)—are better modeled by distinct cognitive mechanisms: attention with syntactic and token memory representations, respectively. This pattern supports our earlier argument regarding the complementary nature of these models (Section 4.2).
@@ 5.1 Do TG’s advantages stem from the COMPOSE attention?
| TG | 96.1 | *_nae | ∗∗∗ (3/3) |
| ( 15.9) | *_nae_so | ∗∗ (3/3) | |
| TG | 86.3 | *_nae | ∗∗∗ (3/3) |
| ( 31.8) | *_nae_so | n.s. (0/3) |
As described in Section 2.2, TG’s key feature is the COMPOSE attention, which explicitly generates single vector representations for closed phrases. Here, we investigate whether TG’s predictive power derives from merely considering syntactic structures or from explicitly treating closed phrases as single representations (see Hale et al., 2018; Brennan et al., 2020). To address this question, we developed TG, a TG variant that processes each action in the action sequence as an individual token (i.e., Choe and Charniak’s approach), without the COMPOSE attention. We trained TG with identical hyperparameters as TG. The baseline regression model (Equation 3.2) was augmented with (i) TG’s surprisal and (ii) Transformer’s NAE to (i) ensure a fair comparison between TG and TG and (ii) distinguish between the effects of direct terminal token access and syntactic structure consideration in TG.
Table 2 presents the LogLik values obtained when incorporating either TG’s or TG’s NAE as fixed effects into the baseline regression model. Note that due to the potential multicollinearity between Transformer’s NAE and TG/TG’s NAE, we focus on the LogLik values and significance of the contribution rather than individual effect sizes. Our analysis reveals two key findings. First, TG’s NAE demonstrates significant predictive power for reading times, even in the presence of Transformer’s NAE, implying that consideration of syntactic structures alone captures certain memory retrievals based on syntactic information. Second, although we should note that the standard deviation across seeds is relatively large, TG’s NAE outperforms TG’s, suggesting that the attention mechanism operating on syntactic memory representations more effectively captures variance in syntax-based memory retrieval. Likelihood ratio tests further revealed that TG’s NAE captured reading time patterns unexplainable by TG (‘TG & TG’‘TG’, 0.001). However, we also found that TG, despite its lower overall predictive power, also explained unique variance (‘TG & TG’‘TG’, 0.001).
To investigate the underlying mechanisms of this complementary relationship, we analyzed the RMSE differences across POS tags (Figure 4). The results showed that TG’s NAE demonstrated advantages over TG’s NAE across most POS tags, with notable exceptions for VBG, VBP, and WP—POS tags that typically involve immediate modification of the previously composed phrase (e.g., (NP (NP ... NP) (SBAR (WHNP (WP who). This pattern leads us to hypothesize that memory retrieval based on syntactic cues typically operates on composed representations for efficiency, but in contexts where immediate modification of a composed phrase occurs (e.g., relative clauses), humans may strategically retrieve individual lexical elements within the composed phrases.
@@ 5.2 Does TG’s NAE capture interference effects?
Psycholinguistic research has identified two primary types of memory retrieval costs: interference effects, which NAE aims to capture, and decay effects—the cognitive load associated with accessing elements at greater linear distances (e.g., Gibson, 1998, 2000). Here, we examine whether TG’s NAE genuinely captures interference effects by testing its independence from variables that model memory decay effects. For modeling decay effects, we employed Category Locality Theory (CLT; Isono, 2024),131313Although Dependency Locality Theory (DLT; Gibson, 1998, 2000) is widely recognized as one of the most prominent models for capturing decay effects, we opted for CLT in this study, following Isono’s finding that DLT-based predictors fail to achieve statistical significance in explaining reading times in the Natural Stories corpus. which treats phrases in syntactic structure141414CLT assumes syntactic structure based on Combinatory Categorial Grammar (Steedman, 2000). as representational units of memory and quantifies decay effects using the distance (measured in content words) between an input and the phrases to be composed with it.
To assess independence, we tested whether TG’s NAE and CLT maintain their contributions when simultaneously included in the baseline regression model (Equation 3.2), and examined their independence through likelihood ratio tests.151515As in other likelihood ratio tests, we used surprisal and NAE values averaged across random seeds. The results (Table 3) show that TG’s NAE exhibited significant effects in both immediate and spillover conditions, and CLT demonstrated a significant immediate effect (with a marginally significant spillover effect). A nested model comparison confirmed that these effects were independent (‘TG & CLT’‘CLT’, ; ‘TG & CLT’‘TG’, ).
@@ 6 Level of description
In cognitive modeling studies based on surprisal theory, explanations typically follow the form “if these LMs were models of human prediction, the difficulty of next-word disambiguation that humans solve would be approximated as follows.” Such explanations typically operate at the most abstract of Marr’s three levels of description—the computational level. Recently, Futrell et al. (2020) proposed lossy-context surprisal to integrate memory representation perspectives into surprisal theory. However, as they explicitly stated, this theory remains at the computational level, relaxing assumptions about memory representations in human predictive processing. In contrast, cognitive models of human memory, such as cue-based retrieval, generally provide explanations about mechanisms that deal with specific mental representations. These explanations typically move down one level to the algorithmic level of description. While not explicitly stated by the authors, we argue that the work of Ryu and Lewis and Oh and Schuler—conceptualizing attention mechanisms as implementations of cue-based retrieval—also operates at the algorithmic level, similar to cue-based retrieval itself. Our research can be characterized as an investigation into the nature of memory representations, addressing fundamental questions at this level (see Hale, 2014).
@@ 7 Conclusion
In this paper, we have demonstrated that attention can serve as the general algorithm for modeling human memory retrieval from two representational systems. Furthermore, we have shown that among the LMs examined in this study (TG, TG, and Transformer), TG—whose attention mechanism uniquely operates on syntactic structures as representational units—best captures dominant factors in human sentence processing. Our results suggest that the integration of attention mechanisms (developed in NLP) with syntactic structures (theorized in linguistics) constitutes a broad-coverage candidate implementation for human memory retrieval. We hope these findings will foster greater collaboration between these two fields.
@@
Our NAE calculation comprised three steps: (i) computing NAE for each attention head in the topmost layers, (ii) adding the values across heads, and (iii) summing subword-level values into word level. While this procedure strictly adhered to Oh and Schuler (2022), alternative approaches to handling layers, attention heads (Ryu and Lewis, 2021), and subword tokens (Oh and Schuler, 2024; Giulianelli et al., 2024) warrant investigation.
While our study provides an in-depth investigation using the Natural Stories corpus—an English self-paced reading time dataset—the breadth of our analysis has certain limitations. The generalizability of our findings to different languages (e.g., Japanese self-paced reading time corpus from Asahara, 2022) and other cognitive load (e.g., gaze duration from Kennedy et al., 2003 or EEG and fMRI from Bhattasali et al., 2020) remains to be investigated.
As discussed in Section 3.1, we employed “perfect oracles” as syntactic structures behind token sequences. This idealization leaves the resolution of local ambiguities, which humans encounter during actual sentence processing, outside the scope of our study (for a conceptual case study, see Appendix D). By incorporating these kinds of entirely new factors, more detailed models could emerge.
Following prior work (Wolfman et al., 2024), we adopted the default TG implementation of a top-down parsing strategy. However, psycholinguistic literature has suggested that a left-corner parsing strategy might be more plausible for human sentence processing (Abney and Johnson, 1991; Resnik, 1992). While previous studies have primarily evaluated the plausibility of parsing strategies from a memory capacity perspective (cf. stack_count), TG’s NAE might offer a new opportunity to revisit the question of cognitively plausible parsing strategies from the perspective of memory interference.
Finally, while this paper focused on investigating the attention mechanism of TG through the lens of memory-based theory, exploring TG as an integrated implementation for expectation-based theory (via surprisal) and memory-based theory (via NAE) represents a promising future direction (Michaelov et al., 2021; Ryu and Lewis, 2022). Specifically, future work could investigate the attention mechanism of TG as the underlying driver of surprisal’s predictive power (Appendix E), analyzing the relationship between surprisal and NAE.
@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant Number 24H00087, Grant-in-Aid for JSPS Fellows JP24KJ0800, JST PRESTO Grant Number JPMJPR21C2, and JST SPRING Grant Number JPMJSP2108.
@@ Appendix A Hyperparameters
Appendix B Correlations between predictors
@@ Appendix C Part-of-speech tags
| -LRB- | Left round bracket | , | Comma |
| -RRB- | Right round bracket | . | Period |
@@ Appendix D Parallel parsing experiment
| TG | 76.6 (6.6) | tg_bs_nae | 2.53 ( 0.2) | 0.001 | 3/3 |
| tg_bs_nae_so | 0.849 ( 0.1) | 0.001 | 3/3 | ||
| Transformer | 28.0 (6.9) | tf_nae | 1.58 ( 0.2) | 0.001 | 3/3 |
| tf_nae_so | 0.457 ( 0.2) | 0.006–0.13 | 1/3 |
As a conceptual case study for the local ambiguity resolution in syntactic structures behind token sequences, we implemented TG’s NAE calculation using 10 syntactic structures obtained through word-synchronous beam search (Stern et al., 2017) with Recurrent Neural Network Grammar (RNNG; Dyer et al., 2016; Kuncoro et al., 2017; Noji and Oseki, 2021).171717https://github.com/aistairc/rnng-pytorch181818RNNG was trained on BLLIP-lg using default hyperparameters. For inference, action beam size and fast track were set to 100 and 1, respectively. NAE was computed individually for each syntactic structure and then aggregated as a weighted average:
where represents the set of syntactic structures synchronized at the -th word ().191919stack_count was similarly calculated as the weighted average across syntactic structures in .
The analysis revealed patterns consistent with those observed when considering only the globally correct syntactic structure: both LMs’ NAE demonstrated significant predictive power for reading times, with TG’s NAE showing stronger contributions compared to Transformer’s (Table 6).202020_bs indicates beam search. The likelihood ratio test further confirmed independent contributions from both LMs ( for both comparisons: ‘TG & Transformer’‘Transformer’ and ‘TG & Transformer’‘TG’).
@@ Appendix E Surprisal experiment
| TG | 265 (11) | tg_surp | 4.63 ( 0.1) | 0.001 | 3/3 |
| tg_surp_so | 1.64 ( 0.1) | 0.001 | 3/3 | ||
| Transformer | 299 (30) | tf_surp | 5.31 ( 0.2) | 0.001 | 3/3 |
| tf_surp_so | 1.68 ( 0.2) | 0.001 | 3/3 |
We analyzed each LM’s surprisal contribution to reading time prediction using a baseline regression model that excluded both LMs’ surprisal from Equation 3.2 but included their NAE (Table 7). While both LMs’ surprisal demonstrated significant predictive power for reading times, Transformer’s surprisal exhibited a stronger contribution compared to TG’s. Additionally, our likelihood ratio test using the averaged surprisal revealed that the regression model incorporating both LMs’ surprisal showed significantly higher predictive power compared to models with only one LM ( for both comparisons: ‘TG & Transformer’‘Transformer’ and ‘TG & Transformer’‘TG’). These findings suggest that (i) unlike attention mechanisms, next-word prediction based solely on token sequences more effectively captures dominant factors of human prediction processing, but (ii) similar to attention mechanisms, both types of next-word prediction—those based on token sequences alone and those leveraging both syntactic structures and token sequences—may coexist as models that capture distinct aspects of human predictive processing.
Appendix F Comparision between TG and TG with a weak baseline regression model

### v2

@@ 4.2 Do TG’s and Transformer’s NAE have independent contributions?
4.3 What aspects of memory retrieval do TG’s and Transformer’s NAE capture?
To investigate the aspects of human memory retrieval captured by TG’s and Transformer’s NAE, we analyzed differences in prediction improvement across part-of-speech (POS) tags annotated in the Natural Stories corpus.141414For a complete list of POS tags in the Natural Stories corpus, see Appendix C. Our analysis followed three steps: (i) selecting POS tags with more than 1,000 occurrences, (ii) for each POS tag, testing the significance of improvement from the baseline regression model (measured in Root Mean Squared Error, RMSE) when adding NAE of the current and previous word as fixed effects,151515We used the same regression models as in Section 4.2, where surprisal and NAE values were averaged across seeds. and (iii) examining the significance of differences in RMSE between TG and Transformer for POS tags where either model showed significant improvement. We assessed significance using Wilcoxon signed-rank tests with Bonferroni correction ().
Figure 3 presents the differences in prediction improvement across POS tags. Consistent with the larger LogLik value, TG’s NAE demonstrated advantages over Transformer’s NAE across a broader range of POS tags. Notably, TG’s NAE exhibited superior improvement across verbs (VB, VBG, VBN, and VBP), while Transformer’s NAE excelled for nouns (NN and NNP). This pattern aligns with our earlier argument regarding the complementary nature of these models (Section 4.2), indicating that different types of retrieval operations—verb-triggered retrieval, which often relies on syntactic features (e.g., argument structure), and noun-triggered retrieval, which often relies on semantic features (e.g., referential associations)—are better captured by distinct attention mechanisms: attention with syntactic and token memory representations, respectively.
@@ 5.1 Do TG’s advantages stem from the COMPOSE attention?
TG | 46.1 | *_nae | ∗∗ (2/3) |
( 9.1) | *_nae_so | ∗∗∗ (3/3) | |
TG | 18.1 | *_nae | ∗∗ (1/3) |
( 9.3) | *_nae_so | ∗∗∗ (3/3) |
As described in Section 2.3, TG’s key feature is the COMPOSE attention, which explicitly generates single vector representations for closed phrases. Here, we investigate whether TG’s predictive power derives from merely considering syntactic structures or from explicitly treating closed phrases as single representations (see Hale et al., 2018; Brennan et al., 2020). To address this question, we developed TG, a TG variant that processes each action in the action sequence as an individual token without the COMPOSE attention (i.e., Choe and Charniak’s Parsing as Language Modeling approach). We trained TG with identical hyperparameters as TG. The baseline regression model (Equation 3.2) was augmented with (i) TG’s surprisal and (ii) Transformer’s NAE to (i) ensure a fair comparison between TG and TG and (ii) distinguish between the effects of direct terminal token access and syntactic structure consideration in TG.161616For direct comparison between TG and TG under the baseline regression model without Transformer’s NAE, see Appendix F.
Table 2 presents the LogLik values obtained when incorporating either TG’s or TG’s NAE as fixed effects into the baseline regression model. Note that due to the potential multicollinearity between Transformer’s NAE and TG/TG’s NAE, we focus on the LogLik values and significance of the contribution rather than individual effect sizes.
Our analysis reveals two key findings. First, TG’s NAE demonstrates significant predictive power for reading times, even in the presence of Transformer’s NAE, implying that consideration of syntactic structures alone captures certain memory retrievals that the token-level attention mechanism cannot capture. Second, TG’s NAE outperforms TG’s, suggesting that the attention mechanism that treats closed phrases as single representations more effectively captures variance in syntax-based memory retrieval. The likelihood ratio tests further revealed that TG’s NAE captured reading time patterns unexplainable by TG (‘TG & TG’‘TG’, ), while TG did not explain unique variance beyond what TG already accounts for (‘TG & TG’‘TG’, ).
We analyzed the RMSE differences across POS tags to investigate which aspects of human memory retrieval are better captured by the COMPOSE attention. Figure 4 showed that TG’s NAE was consistently superior to TG’s NAE across verbs (VB, VBG, and VBN), highlighting the critical role of the COMPOSE attention in capturing verb-triggered retrieval, a type of retrieval that was identified as distinctively better captured by TG compared to vanilla Transformers (Section 4.3).
@@ 5.2 Does TG’s NAE capture interference effects?
Psycholinguistic research has considered two primary types of memory retrieval costs: interference effects, which NAE aims to capture, and decay effects—the cognitive load associated with accessing elements at greater linear distances (e.g., Gibson, 1998, 2000). Here, we examine whether TG’s NAE genuinely captures interference effects by testing its independence from variables that model memory decay effects. For modeling decay effects, we employed Category Locality Theory (CLT; Isono, 2024),171717Although Dependency Locality Theory (DLT; Gibson, 1998, 2000) is widely recognized as one of the most prominent models for capturing decay effects, we opted for CLT here, following Isono’s finding that DLT-based predictors fail to achieve statistical significance in explaining reading times in the Natural Stories corpus. which treats phrases in syntactic structure181818CLT assumes syntactic structure based on Combinatory Categorial Grammar (Steedman, 2000). as representational units of memory and quantifies decay effects using the distance (measured in content words) between an input and the phrases to be composed with it.
To assess independence, we tested whether TG’s NAE and CLT maintain their contributions when simultaneously included in the baseline regression model (Equation 3.2), and examined their independence through likelihood ratio tests.191919As in other likelihood ratio tests, we used surprisal and NAE values averaged across random seeds. The results (Table 3) show that TG’s NAE exhibited significant effects in both immediate and spillover conditions, and CLT demonstrated a significant spillover effect. The likelihood ratio tests confirmed that these effects were independent (‘TG & CLT’‘CLT’, ; ‘TG & CLT’‘TG’, ).
@@ 6 Level of description
In cognitive modeling studies based on surprisal theory, explanations typically follow the form “if these LMs were models of human prediction, the difficulty of next-word disambiguation that humans solve would be approximated as follows.” Such explanations typically operate at the most abstract of Marr’s three levels of description—the computational level. Recently, Futrell et al. (2020) proposed lossy-context surprisal to integrate memory representation perspectives into surprisal theory. However, as they explicitly stated, this theory remains at the computational level, relaxing assumptions about memory representations in human predictive processing. In contrast, cognitive models of human memory, such as cue-based retrieval, generally provide explanations about mechanisms that deal with specific mental representations. These explanations typically move down one level to the algorithmic level of description.
While not explicitly stated by the authors, we argue that the work of Ryu and Lewis (2021) and Oh and Schuler (2022)—conceptualizing attention mechanisms as implementations of cue-based retrieval—represents movement toward the algorithmic level, approaching cue-based retrieval itself. Our research advances this direction by investigating a fundamental question at this level: the nature of memory representations (see Hale, 2014). Future work could fully operate at the algorithmic level by incorporating more realistic elements such as parallel parsing, left-corner parsing strategies, and memory decay mechanisms.
@@ 7 Conclusion
In this paper, we have demonstrated that attention can serve as the general algorithm for modeling human memory retrieval from two representational systems. Furthermore, we have shown that among the LMs examined in this paper (TG, TG, and Transformer), TG—whose attention mechanism uniquely operates on syntactic structures as representational units—best captures dominant factors in human sentence processing. Our results suggest that the integration of attention mechanisms (developed in NLP) with syntactic structures (theorized in linguistics) constitutes a broad-coverage candidate implementation for human memory retrieval. We hope these findings will foster greater collaboration between these two fields.
@@
Our NAE calculation comprised three steps: (i) computing NAE for each attention head in the top layer, (ii) adding the values across heads, and (iii) summing subword-level values into word level. While this procedure strictly adhered to Oh and Schuler (2022), alternative approaches to handling layers, attention heads (Ryu and Lewis, 2021), and subword tokens (Oh and Schuler, 2024; Giulianelli et al., 2024) warrant investigation.
While our study provides an in-depth investigation using the Natural Stories corpus—an English self-paced reading time corpus containing many interesting syntactic constructions—the breadth of our analysis has certain limitations. The generalizability of our findings to different languages (e.g., Japanese self-paced reading time corpus from Asahara, 2022) and other cognitive load (e.g., gaze duration from Kennedy et al., 2003 or EEG and fMRI from Bhattasali et al., 2020) remains to be investigated.
As discussed in Section 3.1, we employed “perfect oracles” as syntactic structures behind token sequences. While this assumption has been widely adopted in previous studies on human syntactic processing (Brennan, 2016; Shain et al., 2016; Stanojević et al., 2021; Isono, 2024), this idealization leaves the resolution of local ambiguities, which humans encounter during actual sentence processing, outside the scope of our study (for a conceptual case study, see Appendix D).
Although we adopted the default TG implementation of a top-down parsing strategy, psycholinguistic literature has suggested that a left-corner parsing strategy might be more plausible for human sentence processing from a perspective of memory capacity (cf. stack_count) (Abney and Johnson, 1991; Resnik, 1992). However, when contrasting memory representations based on syntactic structures versus token sequences, the attention mechanism of the top-down TG can serve as a reasonable approximation of retrieval from structure-based memory—this aligns with previous work contrasting predictions based on syntactic structures versus token sequences, which used top-down TG or RNNG to represent structure-based prediction (Wolfman et al., 2024; Hale et al., 2018; Brennan et al., 2020).
Finally, while this paper focused on investigating the attention mechanism through the lens of memory-based theory, exploring TG and Transformer as integrated implementations for expectation-based theory (via surprisal) and memory-based theory (via NAE) represents a promising future direction (Michaelov et al., 2021; Ryu and Lewis, 2022). Specifically, future work could investigate the attention mechanism as the underlying driver of surprisal’s predictive power (Appendix E), analyzing the relationship between surprisal and NAE.
@@ Acknowledgments
We sincerely thank Michael Wolfman and John Hale for providing detailed information on Wolfman et al. (2024). We also appreciate the insightful reviews provided by anonymous ARR 2025 February reviewers. This work was supported by JSPS KAKENHI Grant Number 24H00087, Grant-in-Aid for JSPS Fellows JP24KJ0800, JST PRESTO Grant Number JPMJPR21C2, and JST SPRING Grant Number JPMJSP2108.
@@ Appendix A Hyperparameters
Appendix B Correlations between the predictors
The correlations between the predictors in our statistical analysis are shown in Table 5. While the NAE from different LMs shows very high correlations with each other, TG and Transformer provide independent predictive power for the self-paced reading times; TG subsumes the predictive power of TG (see Section 4.2 and 5.1).202020_mcomp indicates .
@@ Appendix C Part-of-speech tags
-LRB- | Left round bracket | , | Comma |
-RRB- | Right round bracket | . | Period |
@@ Appendix D Parallel parsing experiment
TG | 56.0 (5.5) | tg_bs_nae | 1.08 ( 0.1) | 0.001 | 3/3 |
tg_bs_nae_so | 2.01 ( 0.1) | 0.001 | 3/3 | ||
Transformer | 26.1 (8.1) | tf_nae | 0.95 ( 0.1) | 0.001 | 3/3 |
tf_nae_so | 1.25 ( 0.2) | 0.001 | 3/3 |
As a conceptual case study for the local ambiguity resolution in syntactic structures behind token sequences, we implemented TG’s NAE calculation using 10 syntactic structures obtained through word-synchronous beam search (Stern et al., 2017) with Recurrent Neural Network Grammar (RNNG; Dyer et al., 2016; Kuncoro et al., 2017; Noji and Oseki, 2021).212121https://github.com/aistairc/rnng-pytorch222222RNNG was trained on BLLIP-lg using default hyperparameters. For inference, action beam size and fast track were set to 100 and 1, respectively. NAE was computed individually for each syntactic structure and then aggregated as a weighted average:
where represents the set of syntactic structures synchronized at the -th word (), and denotes the token position corresponding to the -th word in structure .232323stack_count was similarly calculated as the weighted average across syntactic structures in .
The analysis revealed patterns consistent with those observed when considering only the globally correct syntactic structure: both LMs’ NAE demonstrated significant predictive power for reading times, with TG’s NAE showing stronger contributions compared to Transformer’s (Table 6).242424_bs indicates beam search. The likelihood ratio tests further confirmed independent contributions from both LMs ( for both comparisons: ‘TG & Transformer’‘Transformer’ and ‘TG & Transformer’‘TG’).
@@ Appendix E Surprisal experiment
TG | 62.6 (6.8) | tg_surp | 1.36 ( 0.0) | 0.001 | 3/3 |
tg_surp_so | 1.88 ( 0.1) | 0.001 | 3/3 | ||
Transformer | 159 (12) | tf_surp | 2.47 ( 0.1) | 0.001 | 3/3 |
tf_surp_so | 2.98 ( 0.2) | 0.001 | 3/3 |
We analyzed each LM’s surprisal contribution to reading time prediction using a baseline regression model that excluded both LMs’ surprisal from Equation 3.2 but included their NAE (Table 7). While both LMs’ surprisal demonstrated significant predictive power for reading times, Transformer’s surprisal exhibited a stronger contribution compared to TG’s. Additionally, our likelihood ratio tests using the averaged surprisal revealed that the regression model incorporating both LMs’ surprisal showed significantly higher predictive power compared to models with either LM’s surprisal alone ( for both comparisons: ‘TG & Transformer’‘Transformer’ and ‘TG & Transformer’‘TG’). These findings suggest that (i) unlike attention mechanisms, next-token prediction based solely on token sequences more effectively captures dominant factors of human predictive processing, but (ii) similar to attention mechanisms, both types of next-token prediction—those based on token sequences alone and those leveraging both syntactic structures and token sequences—may coexist as models that capture distinct aspects of human predictive processing.
Appendix F Direct comparison of TG and TG